OpenStudy (cooltowl):

Disprove the given non-identity by giving a counter-example, Square root of sin^2 x-9 = sin x-3

OpenStudy (anonymous):

\[\sqrt{\sin^2x-9}=\sin x-3~~?\]

OpenStudy (cooltowl):

Yes that's the equation

OpenStudy (anonymous):

Consider the fact that the right side, under the square root, can be factored: \[\sqrt{(\sin x+3)(\sin x-3)}\] Is this equal to \(\sin x-3\)?

OpenStudy (anonymous):

*left side, not right

OpenStudy (anonymous):

i think you meant sqrt[ sin^2(x-3) ]

OpenStudy (anonymous):

consider also that the left hand side is an imaginary number, since \(\sin^2(x)-9\leq -8\)

OpenStudy (anonymous):

Notice: \[\large\begin{align*}\sqrt{(\sin x+3)(\sin x-3)}&\overbrace{=}^{?}\sin x-3\\ \sqrt{(\sin x+3)(\sin x-3)}&\overbrace{=}^{?}\sqrt{(\sin x-3)(\sin x-3)}\\ \sqrt{(\sin x+3)}\sqrt{(\sin x-3)}&\overbrace{=}^{?}\sqrt{(\sin x-3)}\sqrt{(\sin x-3)}\\ \sqrt{\sin x+3}&\overbrace{=}^{?}\sqrt{\sin x-3}\\ 3&~\not=-3\end{align*}\] @satellite73 is quite right about what s/he said, and this is exactly why. No such number \(x\) will satisfy the equation.

OpenStudy (cooltowl):

Ok thank you so much very helpful

OpenStudy (anonymous):

yw